OpenStudy (anonymous):

If the point (-1/2, y) lies on the line whose equation is 2x - 3y = 6, what is the value of y?

OpenStudy (whpalmer4):

We know that \[2x-3y=6\]we just need to find the value of \(y\) when \(x = -\frac{1}2\) \[2(-\frac{1}{2})-3y=6\]Solve for \(y\)

OpenStudy (anonymous):

-7/3?